A charged capacitor is connected to a resistor and switch all in series. The circuit has a time constant of 1.50 s. Soon after the switch is closed, the charge of the capacitor is 75.0% of its initial charge.
(i) Find the time interval required for the capacitor to reach this charge.
(ii) If R = 250 Kiloohm, what is the value of C?
1
Expert's answer
2021-02-15T17:43:51-0500

(i)

"Q=Q_0e^{-t/\tau},""0.75Q_0=Q_0e^{-t/\tau}""ln(0.75)=ln(e^{-t/\tau}),""ln(0.75)=-\dfrac{t}{\tau},""t=-\tau ln(0.75)=-1.5\ s\cdot ln(0.75)=0.43\ s."

(ii) From the definition of the time constant, we have:


"\tau=RC,""C=\dfrac{\tau}{R}=\dfrac{1.5\ s}{250\cdot10^3\ \Omega}=6.0\cdot10^6\ F=6.0\ \mu F."